a rhombus has a perimeter of 84 cm and an area of 252 cm square.find the lenght of the altitude
Answer by nerdybill(7384)

a rhombus has a perimeter of 84 cm and an area of 252 cm square.find the lenght of the altitude
.
Since all sides are equal, the length of one side is:
84/4 = 21 cm
.
area = (1/2)ba
252 = (1/2)(21)a
504 = (21)a
504/21 = a
24 cm = a (altitude)